# Markov kernels in Mathlib’s probability library

### v1

@@ 1 Introduction
A Markov kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ is a function from a measurable space $\mathcal{X}$ to the set of probability measures on a measurable space $\mathcal{Y}$ satisfying a measurability condition. It can be seen as describing the effect of a random transformation. In particular, a Markov kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ from a probability space maps the probability measure $P$ on $\mathcal{X}$ to a probability measure $\kappa\circ_{m}P$ on $\mathcal{Y}$. One can consider a category with measurable spaces as objects and Markov kernels as morphisms. In that kernel point of view, there isn’t a single measurable space on which every random variable is defined, and we focus instead on the transformations between them.
The object of this paper is to present the implementation, properties and use of Markov kernels in the probability folder of Mathlib [mat20], the mathematical library of the Lean programming language and theorem prover [Ebn+17, MU21]. Markov kernels play a more important role in that library than in most mathematical texts.
@@
We first introduce Markov kernels, detail particular kernels and describe the different ways in which they can be composed in Section 2. We then present the formalization of the disintegration theorem for kernels and its use to define conditional distributions (Section 3). Having explained the main properties of kernels, we then present two examples of their use in Mathlib, first to define conditional independence and independence (Section 4), then to define sub-Gaussian random variables (Section 5). We finally discuss an application of kernels for entropy and Kullback-Leibler divergence (Section 6).
@@
The definitions and theorems presented in this paper were contributed by the author to the probability folder of Mathlib, except for the results of Section 6 on entropy and divergences. The development of kernels, their properties and the disintegration theorem consist of around 10000 lines of code in the Probability/Kernel folder (excluding notably the Ionescu-Tulcea subfolder that was not added by the author of this paper). Results about independence can be found in the Probability/Independence folder (around 4000 lines of code), and the definition of sub-Gaussian random variables is in Probability/Moments (around 2500 lines together with properties of moment generating functions).
@@
This paper contains links to definitions and theorems from the Mathlib github repository, like for example ProbabilityTheory.Kernel. They all point to a recent commit at the time of writing.
@@
The results discussed in Section 6 about entropy were contributed by the author to the PFR project [And+23]. This is a collaborative formalization of the article [Gow+25], in which the authors prove the polynomial Freiman-Ruzsa conjecture of Katalin Marton. A few remarks in Section 6 on Kullback-Leibler and Rényi divergences come from the work of the author and Lorenzo Luccioli on a formalization of information-theoretic divergences in Lean [LD24]. We are in the process of integrating results about divergences into Mathlib: at the time of writing, the definition of Kullback-Leibler divergence is in Mathlib but only a few properties have been ported. Entropy and other divergences have not yet been the object of pull requests.
Markov kernels in other proof assistants
The formalization of Markov kernels in other proof assistants is mostly motivated by the semantics of probabilistic programming languages [Sta20], while our goal is to provide a general-purpose theory of kernels for probability theory. In Isabelle/HOL, [HMS23] formalized quasi-Borel spaces and they defined Markov kernels, finite and s-finite kernels. In Rocq/Coq, [ACS25] also formalized Markov, finite and s-finite kernels, and additionally $\sigma$-finite kernels. They discuss in details the problem of proving the measurability needed to define the composition of kernels, which we will omit in this paper. Both of those formalizations build classes of kernels and some of their compositions (corresponding to what we present in Section 2) but don’t discuss disintegration of kernels or conditional distributions.
Notation
@@
$\mathbb{R}_{+,\infty}$ is the set of nonnegative extended real numbers. $\Omega,\mathcal{T},\mathcal{U},\mathcal{X},\mathcal{Y},\mathcal{Z}$ denote measurable spaces. The set of sets of $\mathcal{X}$ is denoted by $\mathcal{P}(\mathcal{X})$. The set of nonnegative measures on a measurable space $\mathcal{X}$ is written $\mathcal{M}(\mathcal{X})$. A kernel from $\mathcal{X}$ to $\mathcal{Y}$ is denoted by $\mathcal{X}\rightsquigarrow\mathcal{Y}$. The expectation of a random variable $X$ with respect to a measure $\mu$ is written $\mu[X]=\int_{\omega}X(\omega)\,d\mu$. This is a Lebesgue integral if $X$ is takes values in $\mathbb{R}_{+,\infty}$, and a Bochner integral in other cases (see [Gou22] for a discussion of integrals in Mathlib). Other notation, notably about constructions of kernels, will be introduced with the corresponding definitions.
@@ 2 Probability spaces and Markov kernels
We can define a measurable space on the set of measures on $\mathcal{Y}$ as the coarsest $\sigma$-algebra such that the function $\mathcal{M}(\mathcal{Y})\to\mathbb{R}_{+,\infty}$ defined by $\mu\mapsto\mu(B)$ is measurable for every measurable set $B$ of $\mathcal{Y}$. That measurable space was already implemented in Mathlib by Johannes Hölzl in 2019 in the context of the implementation of the Giry monad (see [Höl17] for his work on the same topic in Isabelle/HOL). A kernel is then simply a measurable function from $\mathcal{X}$ to the set of measures on $\mathcal{Y}$. In that sense, kernels were already defined by Hötzl in 2019, and the composition of a measure and a kernel (see Section 2.2) was implemented as the monadic bind of the Giry monad. But apart from a short file containing the definition and that composition, kernels were not used prior to our work.
@@
Kernels are fully specified by their action on measurable functions. That is, if two kernels $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ are such that for every measurable function $f:\mathcal{Y}\to\mathbb{R}_{+,\infty}$ and every $x\in\mathcal{X}$, $\int_{y}f(y)d\kappa(x)=\int_{y}f(y)d\eta(x)$, then $\kappa=\eta$. We will use that fact repeatedly in the following in order to define kernels by giving their integrals.
@@ 2.1 Classes of kernels
A kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ is a Markov kernel if $\kappa(x)$ is a probability measure for every $x\in\mathcal{X}$. If the supremum of $\kappa(x)(\mathcal{Y})$ over $x\in\mathcal{X}$ is finite, then $\kappa$ is said to be a finite kernel. Finally, we also defined the class of s-finite kernels, which are kernels that can be expressed as a countable sum of finite kernels. Those three properties are denoted by typeclasses [IsMarkovKernel κ], [IsFiniteKernel κ] and [IsSFiniteKernel κ] respectively. [VO18] is a good reference on classes of kernels, in particular the less often used s-finite class.
There is a fourth class of kernels that is mentioned in the literature but that we did not implement in Mathlib: $\sigma$-finite kernels. Those are sums of countably many finite kernels that are pairwise singular. They don’t enjoy the nice composition properties of the other classes (see the next subsection), and we decided to wait for a concrete use case before implementing them. Although, as we will discuss further down, the Radon-Nikodym theorem for kernels stated below could be a potential use case.
A measure can be seen as a kernel from the one-point space $\mathbf{1}$ to the space on which it is defined. A random variable $X:\mathcal{X}\to\mathcal{Y}$ can be seen as a kernel $D_{X}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ that maps $x$ to the Dirac measure at $X(x)$. Such a kernel is called a deterministic kernel and is a Markov kernel. Among the deterministic kernels, a few play a special role. The identity kernel $\mathrm{id}_{\mathcal{X}}:\mathcal{X}\rightsquigarrow\mathcal{X}$ is the deterministic kernel associated with the identity function: it maps $x$ to the Dirac measure at $x$. The copy kernel $\Delta_{\mathcal{X}}:\mathcal{X}\rightsquigarrow\mathcal{X}\times\mathcal{X}$ maps $x$ to the Dirac measure at $(x,x)$. The discard kernel $_{\mathcal{X}}:\mathcal{X}\rightsquigarrow\mathbf{1}$ maps $x$ to the unique probability measure on the one-point space $\mathbf{1}$. Another useful (not deterministic) kernel is the constant kernel $\kappa_{\mu}:\Omega\rightsquigarrow\mathcal{X}$ which maps every point of $\Omega$ to the same probability measure $\mu$ on $\mathcal{X}$.
@@ 2.2 Composition of kernels
We use the notation $\circ_{k}$ to reflect the notation in Mathlib, in which we can’t use simply since this is notation for function composition: a kernel $\mathcal{X}\rightsquigarrow\mathcal{Y}$ is a function, but from $\mathcal{X}$ to $\mathcal{M}(\mathcal{Y})$, and composition of kernels is not the same as their composition as functions. We omit the proof that $\eta\circ_{k}\kappa$ is indeed a well defined kernel and refer the reader to [ACS25]. We will also omit similar proofs for other constructions of kernels. A composition of Markov kernels is a Markov kernel, and finite and s-finite kernels are also stable by composition. $\sigma$-finite kernels (which we did not implement) are not stable by composition. For finite kernels, the existence of a uniform bound $C<\infty$ such that $\kappa(x)(\mathcal{Y})<C$ for all $x$ is important to have stability: if we require only that $\kappa(x)(\mathcal{Y})<\infty$ for all $x$, then the composition of two such kernels would be $\sigma$-finite but not necessarily finite. The stability of Markov, finite and s-finite kernels will also be true for the other constructions presented below. For $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Z}$, we write $\kappa_{\mathrm{fst}}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ and $\kappa_{\mathrm{snd}}:\mathcal{X}\rightsquigarrow\mathcal{Z}$ for the compositions of $\kappa$ with the deterministic kernels associated with the first and second projections respectively.
@@
where $\kappa(x)\otimes\eta(t)$ is the product measure of $\kappa(x)$ and $\eta(t)$. Note that here we need the kernels to be s-finite to ensure that the product measure is a measurable function. In Mathlib, the parallel composition is defined for all kernels, but it is set to the default value 0 if either kernel is not s-finite. This is a standard trick in formalization to avoid having to carry hypotheses in every expression involving that parallel composition (they will instead be present as hypotheses of the lemmas) [Buz20]. All kernel constructions that require s-finiteness use the same trick.
@@
Composition of kernels is associative, $\xi\circ_{k}(\eta\circ_{k}\kappa)=(\xi\circ_{k}\eta)\circ_{k}\kappa$, and the name “identity kernel” is justified by the fact that it is a neutral element for that composition, $\mathrm{id}\circ_{k}\kappa=\kappa$ and $\kappa\circ_{k}\mathrm{id}=\kappa$. The composition-product is, mathematically speaking, associative as well, but here the type theory used in Lean makes it more complex to express that property. Indeed, suppose that we have three s-finite kernels $\kappa:\mathcal{T}\rightsquigarrow\mathcal{X}$, $\eta:\mathcal{T}\times\mathcal{X}\rightsquigarrow\mathcal{Y}$ and $\xi:\mathcal{T}\times(\mathcal{X}\times\mathcal{Y})\rightsquigarrow\mathcal{Z}$. Then we can form a kernel $(\kappa\otimes_{k}\eta)\otimes_{k}\xi$ without issue. Associativity would be the statement that this is the same kernel as $\kappa\otimes_{k}(\eta\otimes_{k}\xi)$. However, there are problems with that second kernel: we can’t build $\eta\otimes_{k}\xi$ since the type of $\xi$ is $\mathcal{T}\times(\mathcal{X}\times\mathcal{Y})\rightsquigarrow\mathcal{Z}$ and not $(\mathcal{T}\times\mathcal{X})\times\mathcal{Y}\rightsquigarrow\mathcal{Z}$. The product of types is not associative. Similarly, even if we could obtain $\kappa\otimes_{k}(\eta\otimes_{k}\xi)$, its type would be $\mathcal{T}\rightsquigarrow\mathcal{X}\times(\mathcal{Y}\times\mathcal{Z})$ while the type of $(\kappa\otimes_{k}\eta)\otimes_{k}\xi$ is $\mathcal{T}\rightsquigarrow(\mathcal{X}\times\mathcal{Y})\times\mathcal{Z}$. We can insert explicitly compositions with associators to fix the type issues, but that makes the statement of associativity more cumbersome.
@@
Since measures can be seen as kernels from the one-point space, we can use these operations to combine measures and kernels. This leads to a composition definition $\kappa\circ_{m}\mu$ for a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ and a measure $\mu$ on $\mathcal{X}$, which is a measure on $\mathcal{Y}$. Similarly, we can define a composition-product $\kappa\otimes_{m}\mu$ for a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ and a measure $\mu$ on $\mathcal{X}$, which is a measure on $\mathcal{X}\times\mathcal{Y}$ (Measure.compProd). For a deterministic kernel $D_{X}$ associated with a random variable $X:\mathcal{X}\to\mathcal{Y}$, $D_{X}\circ_{m}\mu$ is the pushforward measure of $\mu$ by $X$, which we denote by $X_{*}\mu$ and is such that $X_{*}\mu(B)=\mu(X^{-1}(B))$ for all measurable sets $B\subseteq\mathcal{Y}$.
@@
The composition is related to the composition-product by the formula $\eta\circ_{k}\kappa=(\kappa\otimes_{k}\eta)_{\mathrm{snd}}$. In a first attempt at implementing kernels in Mathlib, we defined the composition-product as the primitive operation on kernels, and derived the composition and the product from it. This had the advantage that once the work is done to prove that the composition-product is indeed a kernel, the other operations are easy to define. There is however a drawback: the composition-product requires s-finite kernels, while the composition should be defined for all kernels. Defining the composition as $(\kappa\otimes_{k}\eta)_{\mathrm{snd}}$ is a loss in generality. The change to use composition and parallel composition as basic operations also aligns better with the categorical viewpoint on kernels.

### v2

@@ 1 Introduction
A Markov kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ is a function from a measurable space $\mathcal{X}$ to the set of probability measures on a measurable space $\mathcal{Y}$ satisfying a measurability condition. It can be seen as describing the effect of a random transformation. In particular, a Markov kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ from a probability space maps the probability measure $P$ on $\mathcal{X}$ to a probability measure $\kappa\circ_{m}P$ on $\mathcal{Y}$ (the precise definition of that measure will be given in Section 2.2). One can consider a category with measurable spaces as objects and Markov kernels as morphisms. In that kernel point of view, there is not a single measurable space on which every random variable is defined, and we focus instead on the transformations between them.
The object of this paper is to present the implementation, properties and use of Markov kernels in the probability folder of Mathlib [mathlib], the mathematical library of the Lean programming language and theorem prover [lean, moura2021lean]. Markov kernels play a more important role in that library than in most mathematical texts: they provide a unified way to formalize various concepts of probability theory.
@@
We first introduce Markov kernels, detail particular kernels and describe the different ways in which they can be composed in Section 2. We then present the formalization of the disintegration theorem for kernels and its use to define conditional distributions (Section 3). Having explained the main properties of kernels, we then present two examples of their use in Mathlib, first to define conditional independence and independence (Section 4), then to define sub-Gaussian random variables (Section 5). We finally discuss an application of kernels to entropy and Kullback-Leibler divergence (Section 6).
@@
The definitions and theorems presented in this paper were contributed by the author to the probability folder of Mathlib, except for the results of Section 6 on entropy and divergences that describe work by the author in other repositories. The development of kernels, their properties and the disintegration theorem consist of around 10000 lines of code in the Probability/Kernel folder (excluding notably the Ionescu-Tulcea subfolder that was not added by the author of this paper). Results about independence can be found in the Probability/Independence folder (around 4000 lines of code), and the definition of sub-Gaussian random variables is in Probability/Moments (around 2500 lines together with properties of moment generating functions).
@@
This paper contains links to definitions and theorems from the Mathlib github repository, like for example ProbabilityTheory.Kernel. They all point to a recent commit at the time of writing 111commit dd465758525e4dcf29027e0ebc5d8fd552be361f. The name ProbabilityTheory.Kernel refers to the name of the declaration, not to its path: ProbabilityTheory is a namespace in which we added a Kernel definition, and that declaration ProbabilityTheory.Kernel is placed in the file Probability/Kernel/Defs.lean but could be moved to another file without changing its name.
@@
The results discussed in Section 6 about entropy were contributed by the author to the PFR project [Anderson_Formalization_of_the_2023]. This is a collaborative formalization of the article [gowers2025conjecture], in which the authors prove the polynomial Freiman-Ruzsa conjecture of Katalin Marton. A few remarks in Section 6 on Kullback-Leibler and Rényi divergences come from the work of the author and Lorenzo Luccioli on a formalization of information-theoretic divergences in Lean [TestingLowerBounds]. We are in the process of integrating results about divergences into Mathlib: at the time of writing, the definition of Kullback-Leibler divergence is in Mathlib but only a few properties have been ported. Entropy and other divergences have not yet been the object of pull requests.
Notations
@@
$\mathbb{R}_{+,\infty}$ is the set of nonnegative extended real numbers. $\Omega,\mathcal{T},\mathcal{U},\mathcal{X},\mathcal{Y},\mathcal{Z}$ denote measurable spaces. The set of sets of $\mathcal{X}$ is denoted by $\mathcal{P}(\mathcal{X})$. The set of nonnegative measures on a measurable space $\mathcal{X}$ is written $\mathcal{M}(\mathcal{X})$. A kernel from $\mathcal{X}$ to $\mathcal{Y}$ is denoted by $\mathcal{X}\rightsquigarrow\mathcal{Y}$. The expectation of a random variable $X$ with respect to a measure $\mu$ is written $\mu[X]=\int_{\omega}X(\omega)\,d\mu$. This is a Lebesgue integral if $X$ is takes values in $\mathbb{R}_{+,\infty}$, and a Bochner integral in other cases (see [gouezel2022formalization] for a discussion of integrals in Mathlib). The notation $f:\mathcal{X}\to\mathcal{Y}\to\mathcal{Z}$ means that $f$ is a curried function $\mathcal{X}\to(\mathcal{Y}\to\mathcal{Z})$ and we write its application to $x$ and $y$ as $f(x)(y)$. Other notation, notably about constructions of kernels, will be introduced with the corresponding definitions.
@@ 2 Probability spaces and Markov kernels
We can define a measurable space on the set of measures on $\mathcal{Y}$ as the coarsest $\sigma$-algebra such that the function $\mathcal{M}(\mathcal{Y})\to\mathbb{R}_{+,\infty}$ defined by $\mu\mapsto\mu(B)$ is measurable for every measurable set $B$ of $\mathcal{Y}$. That measurable space was already implemented in Mathlib by Johannes Hölzl in 2019 in the context of the implementation of the Giry monad (see [holzl2017markov] for his work on the same topic in Isabelle/HOL). Mathematically, a kernel is then simply a measurable function from $\mathcal{X}$ to the set of measures on $\mathcal{Y}$. In that sense, kernels were already defined by Hötzl in 2019, and the composition of a measure and a kernel (see Section 2.2) was implemented as the monadic bind of the Giry monad. But apart from a short file containing the definition and that composition, kernels were not used prior to our work.
@@
Kernels are fully specified by their action on measurable functions. That is, if two kernels $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ are such that for every measurable function $f:\mathcal{Y}\to\mathbb{R}_{+,\infty}$ and every $x\in\mathcal{X}$, $\int_{y}f(y)d\kappa(x)=\int_{y}f(y)d\eta(x)$, then $\kappa=\eta$.
theorem ext_fun_iff : κ = η x f, Measurable f ∫⁻ b, f b ∂κ x = ∫⁻ b, f b ∂η x
The notation ∫⁻ indicates that this is a Lebesgue integral.
We will use that fact repeatedly in the following in order to describe kernels by giving their integrals. Another way to show that two kernels are equal is to show that their values $\kappa(x)$ and $\eta(x)$ are equal for all $x$, which since they are measures can be checked by checking that they give the same value to all measurable sets.
@@ 2.1 Classes of kernels
A kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ is a Markov kernel if $\kappa(x)$ is a probability measure for every $x\in\mathcal{X}$. If the supremum of $\kappa(x)(\mathcal{Y})$ over $x\in\mathcal{X}$ is finite, then $\kappa$ is said to be a finite kernel. Finally, we also defined the class of s-finite kernels, which are kernels that can be expressed as a countable sum of finite kernels. Those three properties are denoted by typeclasses [IsMarkovKernel κ], [IsFiniteKernel κ] and [IsSFiniteKernel κ] respectively. [vakar2018s] is a good reference on classes of kernels, in particular the less often used s-finite class.
There is a fourth class of kernels that is mentioned in the literature but that we did not implement in Mathlib: $\sigma$-finite kernels. Those are sums of countably many finite kernels that are pairwise singular. They do not enjoy the nice composition properties of the other classes (see the next subsection), and we decided to wait for a concrete use case before implementing them. Although, as we will discuss further down, the Radon-Nikodym theorem for kernels stated below is a potential use case.
A measure can be seen as a kernel from the one-point space $\mathbf{1}$ to the space on which it is defined. A random variable $X:\mathcal{X}\to\mathcal{Y}$ can be seen as a kernel $D_{X}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ that maps $x$ to the Dirac measure at $X(x)$. Such a kernel is called a deterministic kernel and is a Markov kernel. Among the deterministic kernels, a few play a special role. The identity kernel $\mathrm{id}_{\mathcal{X}}:\mathcal{X}\rightsquigarrow\mathcal{X}$ is the deterministic kernel associated with the identity function: it maps $x$ to the Dirac measure at $x$. The copy kernel $\Delta_{\mathcal{X}}:\mathcal{X}\rightsquigarrow\mathcal{X}\times\mathcal{X}$ maps $x$ to the Dirac measure at $(x,x)$. The discard kernel $_{\mathcal{X}}:\mathcal{X}\rightsquigarrow\mathbf{1}$ maps $x$ to the unique probability measure on the one-point space $\mathbf{1}$. Another useful (not deterministic) kernel is the constant kernel $\Omega\rightsquigarrow\mathcal{X}$ which maps every point of $\Omega$ to the same probability measure $\mu$ on $\mathcal{X}$.
@@ 2.2 Composition of kernels
We use the notation $\circ_{k}$ to reflect the notation in Mathlib, in which we cannot use simply $\circ$ since this is notation for function composition: a kernel $\mathcal{X}\rightsquigarrow\mathcal{Y}$ is a function, but from $\mathcal{X}$ to $\mathcal{M}(\mathcal{Y})$, and composition of kernels is not the same as their composition as functions.
To explain how $\eta\circ_{k}\kappa$ is defined, we need to first introduce the composition of a kernel and a measure, which we denote by $\kappa\circ_{m}\mu$ for a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ and a measure $\mu$ on $\mathcal{X}$. This is defined as the bind in the Giry monad [giry1982categorical], which was implemented prior to our work (Measure.bind). The map or pushforward measure of $\mu$ on $\mathcal{X}$ by a measurable function $X:\mathcal{X}\to\mathcal{Z}$, which we denote by $X_{*}\mu$, is the measure such that $X_{*}\mu(B)=\mu(X^{-1}(B))$ for all measurable sets $B\subseteq\mathcal{Z}$. For $\kappa$ and $\mu$, we can thus define the map of $\mu$ by $\kappa$ (seeing $\kappa$ as a measurable function from $\mathcal{X}$ to $\mathcal{M}(\mathcal{Y})$), the measure $\kappa_{*}\mu$ on $\mathcal{M}(\mathcal{Y})$. We can then define the join in the monad, which maps a measure $\nu$ on $\mathcal{M}(\mathcal{Y})$ to the measure $\nu^{\sharp}$ on $\mathcal{Y}$ defined by $\nu^{\sharp}(B)=\int_{\mu}\mu(B)d\nu$ for all measurable sets $B$ of $\mathcal{Y}$. The bind $\kappa\circ_{m}\mu$ is then defined as the join of the map of $\mu$ by $\kappa$: $\kappa\circ_{m}\mu=(\kappa_{*}\mu)^{\sharp}$. We can check that map, join and bind are measurable. We don’t detail the measurability proof here and refer the reader to [affeldt2025semantics]. We will also omit measurability proofs for other constructions of kernels. The composition $\eta\circ_{k}\kappa$ is then defined as the kernel such that for all $x$, $(\eta\circ_{k}\kappa)(x)=\eta\circ_{m}\kappa(x)$.
[…]
toFun a := a).bind η
measurable := (Measure.measurable_bind η.measurable).comp κ.measurable
A composition of Markov kernels is a Markov kernel, and finite and s-finite kernels are also stable by composition. $\sigma$-finite kernels (which we did not implement) are not stable by composition. For finite kernels, the existence of a uniform bound $C<\infty$ such that $\kappa(x)(\mathcal{Y})<C$ for all $x$ is important to have stability: if we require only that $\kappa(x)(\mathcal{Y})<\infty$ for all $x$, then the composition of two such kernels would be $\sigma$-finite but not necessarily finite. The stability of Markov, finite and s-finite kernels will also be true for the other constructions presented below. For $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Z}$, we write $\kappa_{\mathrm{fst}}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ and $\kappa_{\mathrm{snd}}:\mathcal{X}\rightsquigarrow\mathcal{Z}$ for the compositions of $\kappa$ with the deterministic kernels associated with the first and second projections respectively.
@@
in which $\kappa(x)\otimes\eta(t)$ is the product measure of $\kappa(x)$ and $\eta(t)$. Note that here we need the kernels to be s-finite to ensure that the product measure is a measurable function. In Mathlib, the parallel composition is defined for all kernels, but it is set to the default value 0 if either kernel is not s-finite. This is a standard trick in formalization to avoid having to carry hypotheses in every expression involving that parallel composition (they will instead be present as hypotheses of the lemmas) [defaultValues, defaultValues2]. All kernel constructions that require s-finiteness use the same trick.
@@
$\kappa\circ_{k}\eta$ composition of kernels $\kappa$ and $\eta$
$\kappa\circ_{m}\mu$ composition of a kernel $\kappa$ and a measure $\mu$
$X_{*}\mu$ pushforward of a measure $\mu$ by a random variable $X$
$\kappa\parallel_{k}\eta$ parallel composition of kernels $\kappa$ and $\eta$
$\kappa\otimes_{k}\eta$ composition-product of kernels $\kappa$ and $\eta$
$\mu\otimes_{m}\kappa$ composition-product of a measure $\mu$ and a kernel $\kappa$
$\kappa\times_{k}\eta$ product of kernels $\kappa$ and $\eta$
Table 1: Summary of the notations for kernels and measures.
@@
Composition of kernels is associative, $\xi\circ_{k}(\eta\circ_{k}\kappa)=(\xi\circ_{k}\eta)\circ_{k}\kappa$, and the name “identity kernel” is justified by the fact that it is a neutral element for that composition, $\mathrm{id}\circ_{k}\kappa=\kappa$ and $\kappa\circ_{k}\mathrm{id}=\kappa$. The composition-product is, mathematically speaking, associative as well, but here the type theory used in Lean makes it more complex to express that property. Indeed, suppose that we have three s-finite kernels $\kappa:\mathcal{T}\rightsquigarrow\mathcal{X}$, $\eta:\mathcal{T}\times\mathcal{X}\rightsquigarrow\mathcal{Y}$ and $\xi:\mathcal{T}\times(\mathcal{X}\times\mathcal{Y})\rightsquigarrow\mathcal{Z}$. Then we can form a kernel $(\kappa\otimes_{k}\eta)\otimes_{k}\xi$ without issue. Associativity would be the statement that this is the same kernel as $\kappa\otimes_{k}(\eta\otimes_{k}\xi)$. However, there are problems with that second kernel: we cannot build $\eta\otimes_{k}\xi$ since the type of $\xi$ is $\mathcal{T}\times(\mathcal{X}\times\mathcal{Y})\rightsquigarrow\mathcal{Z}$ and not $(\mathcal{T}\times\mathcal{X})\times\mathcal{Y}\rightsquigarrow\mathcal{Z}$. Those two product of types cannot be used interchangeably. Similarly, even if we could obtain $\kappa\otimes_{k}(\eta\otimes_{k}\xi)$, its type would be $\mathcal{T}\rightsquigarrow\mathcal{X}\times(\mathcal{Y}\times\mathcal{Z})$ while the type of $(\kappa\otimes_{k}\eta)\otimes_{k}\xi$ is $\mathcal{T}\rightsquigarrow(\mathcal{X}\times\mathcal{Y})\times\mathcal{Z}$. We can insert explicitly compositions with associators to fix the type issues, but that makes the statement of associativity more cumbersome.
@@
Since measures can be seen as kernels from the one-point space, we can use these operations to combine measures and kernels. From $\mu$ a measure on $\mathcal{X}$ we can build a kernel $c_{\mu}:\mathbf{1}\rightsquigarrow\mathcal{X}$ that maps the unique point of $\mathbf{1}$ to $\mu$. We can then compose $c_{\mu}$ and $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ to get a kernel $\kappa\circ_{k}c_{\mu}:\mathbf{1}\rightsquigarrow\mathcal{Y}$, which takes a single value and defines a measure on $\mathcal{Y}$. This recovers the composition definition $\kappa\circ_{m}\mu$ for a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ and a measure $\mu$ on $\mathcal{X}$. By similarly seeing measures as kernels from the one-point space, we can define a composition-product $\mu\otimes_{m}\kappa$ for a measure $\mu$ on $\mathcal{X}$ and a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$, which is a measure on $\mathcal{X}\times\mathcal{Y}$ (Measure.compProd). For a deterministic kernel $D_{X}$ associated with a random variable $X:\mathcal{X}\to\mathcal{Y}$, $D_{X}\circ_{m}\mu$ is the pushforward measure of $\mu$ by $X$, the measure $X_{*}\mu$ on $\mathcal{Y}$.
@@
The composition is related to the composition-product by the formula $\eta\circ_{k}\kappa=(\kappa\otimes_{k}\eta)_{\mathrm{snd}}$ (in which abuse notation: we see $\eta:\mathcal{Y}\rightsquigarrow\mathcal{Z}$ as a kernel $\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathcal{Z}$ by ignoring the first component). In a first attempt at implementing kernels in Mathlib, we defined the composition-product as the primitive operation on kernels, and derived the composition and the product from it. This had the advantage that once the work is done to prove that the composition-product is indeed a kernel, the other operations are easy to define. There is however a drawback: the composition-product requires s-finite kernels, while the composition should be defined for all kernels. Defining the composition as $(\kappa\otimes_{k}\eta)_{\mathrm{snd}}$ is a loss in generality. The change to use composition and parallel composition as basic operations also aligns better with the categorical viewpoint on kernels.